Express each number as a product of its prime factors: 
3825

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

3825
By Taking the LCM of 3825, we will get the product of its prime factor.
Hence,3825 = 3 × 3 × 5 × 5 × 17 × 1 = 3^2×5^2×17

Thanks